# Few-Shot Domain Adaptation for Learned Image Compression

### v1

@@ Abstract
Index Terms:
Learned Image Compression, Domain Adaptation, Few-Shot Learning.
I Introduction
With the explosion of images in modern multimedia, image compression has become crucial for efficient storage and transmission. Over the past decade, conventional standards including JPEG [45], H.265/HEVC [41] and H.266/VVC [9], have significantly contributed to the real-world applications of image compression. Recently, learned image compression (LIC) has made remarkable strides in rate-distortion (RD) performance by leveraging non-linear transforms and end-to-end optimization. Several studies [19; 30; 21] have even surpassed VTM (the reference software for H.266/VVC) on both peak signal-to-noise ratio (PSNR) and multi-scale structural similarity (MS-SSIM), indicating the immense potential of LIC for next-generation image compression techniques.
@@ I Introduction
Despite the encouraging progress, some studies [38; 32; 42] have pointed out the performance degradation of pre-trained LIC models on out-of-domain images (we simplify in/out-of-training-domain as in/out-of-domain for convenience). It is common practice to train and evaluate LIC models on natural image benchmarks such as DIV2K [1], Kodak [16], Tecnick [2], and CLIC [12]. However, practical codecs may well be challenged by images from various domains, many of which have significant domain gaps compared to natural images. For demonstration purposes, we test advanced LIC models including Cheng2020 [11], ELIC [19], TCM [30], and MLIC++ [21] in Fig. 1. While all these models perform comparably or better than the conventional codec VTM on natural images, they exhibit varying degrees of performance drop and generally fall behind VTM on out-of-domain images such as pixel-style art and screen content, highlighting the necessity for efficient adaptation methods for pre-trained LIC models.
@@
In this regard, most existing methods [38; 32; 42] suggest instance adaptation (IA), which aims to adapt LIC models to a single image at inference. Despite its effectiveness, IA does not perform actual domain adaptation (DA) for pre-trained models, and is time-consuming and computationally expensive due to its per-image online training. On the other hand, Katakol et al. [24] proposed DA for LIC using limited target samples. They suggested selective finetune, and adapted pre-trained models to a target domain with one-time training and transmission, demonstrating great adaptation efficiency for deployed models. Nevertheless, the proposed method is restricted to GDN structures [3], and suffers from large additional parameters on recent schemes [19; 30; 22].
@@
Despite the promising applications like codec calibration for batch transmission, DA for LIC remains largely unexplored. Recently, Presta et al. [35] pre-trained a gate network on multiple domains, enhancing the generalization capacity of LIC models. However, they did not address the problem of adapting LIC models to a specific domain. Therefore, in this paper, we focus on adapting pre-trained LIC models to a target domain. For practicality, we follow Katakol et al. [24] and study DA for LIC with limited samples. We also highlight the versatility of the DA method to different domains and mainstream LIC schemes.
@@
Building on these insights, we propose a universal few-shot domain adaptation method for LIC using compact adapters, achieving superior performance across various domains and mainstream LIC schemes. Inspired by [28], we explore domain gaps in LIC through disturbed channel-wise decomposition. We demonstrate that pre-trained LIC models suffer from scattered channel-wise energy allocation on out-of-domain images. To address this problem, we introduce convolution-based adapters (Conv-Adapters) and low-rank adapters (LoRA-Adapters) to the pre-trained models for channel-wise re-allocation. These adapters are trained with a few target samples, and are plug-and-play for processing specific domains.
@@
II Related Work
II-A Learned Image Compression
Learned image compression (LIC) has demonstrated competitive potential against traditional codecs. Ballé et al. [4] proposed the first end-to-end LIC model with non-linear transforms and uniform quantizer, which was further strengthened by hyperprior [5] and context model [33]. Cheng et al. [11] first achieved comparable performance with VTM using Gaussian mixture likelihoods. He et al. [19] proposed uneven channel autoregressive model, demonstrating pleasant performance and complexity. Recently, Liu et al. [30] designed Transformer-CNN mixture block to incorporate the ability of both structures for LIC. Jiang et al. [22] introduced multi-reference context to capture different correlations, achieving the state-of-the-art performance [21].
@@ II-A Learned Image Compression
In the regard of interpretability, early work [5; 11] visualized the latent in LIC to evaluate the preciseness of the entropy model. He et al. [19] sorted latent channels by energy, suggesting low-frequency components concentrated on a few channels while the rest was extremely sparse. Duan et al. [14] interpreted latent channels as orthogonal transform coefficients, while Li et al. [28] suggested frequency decomposition is an intrinsic property for LIC. Nevertheless, these results on latent representation did not fully consider out-of-domain images.
@@
II-B Adaptation for Learned Image Compression
To alleviate domain gaps in LIC, two kinds of adaptation namely instance adaptation (IA) and domain adaptation (DA) have been studied for LIC. IA [10; 43; 42; 38; 32] adapts pre-trained models to a single image rather than a domain using per-image online training, including content [49; 10] and decoder [44; 43; 51; 42; 38; 32] adaptation. The additional parameters are compressed and transmitted together with bitstream. Despite its effectiveness, IA is computationally expensive and time-consuming due to its per-image fashion.
Contrary to IA, Katakol et al. [24] introduced the problem of DA for LIC, adapting pre-trained LIC models to a target domain using multiple samples and one-time training. They suggested finetuning GDN layers [3] with channel-wise parameters and the entire entropy model. However, this is restricted by specific network structures, and additional parameters due to the complex entropy models. We alleviate this issue by designing compact and universal adapters.
The scope of this work belongs to DA, which is distinct from IA on several dimensions: (a) Our goal is transferring pre-trained LIC models to a target domain rather than a specific instance. Besides, (b) our method adopts one-time training and transmission rather than the per-image fashion at inference. The additional parameters are evaluated alone and are independent to bitstream and bit-rate. In addition, we suggest this work is complementary to instance adaptation, as we demonstrate the compatibility of both techniques for further RD improvement on out-of-domain images.
II-C Few-Shot Domain Adaptation
Domain adaptation [37; 39] aims to bridge the gap between the source domain and the target domain, which exists widely in deep learning. In practice, collecting sufficient data from a specific domain can be expensive and difficult due to various limitations. Therefore, the combination of few-shot learning [15; 40] and domain adaptation plays a critical role. Adapting on a few target samples, models pre-trained on a source domain are expected to perform similarly on a target domain [27]. Although few-shot domain adaptation has been well explored in many computer vision tasks, like image classification [29; 13], segmentation [25; 31; 23], and generation [17; 26; 50; 46], limited attention has been paid to learned image compression, and there still lack universal few-shot domain adaptation methods for mainstream learned image compression schemes and domains.
III Method
III-A Motivation
Recent studies [14; 28] suggest that the transform in LIC can be interpreted as channel-wise frequency decomposition in a non-linear fashion. We compare latent channels to frequency components of a source image using linear transforms (e.g. FFT), as reconstructions from partial channels or partial frequency components demonstrate obvious similarities in Fig. 2. Motivated by this, we further study the channel-wise decomposition on different domains by reconstructions from specific channels in Fig. 3 (c) and (d). Consistent with previous findings [19; 14; 28], decomposed information concentrates on high-energy channels (respectively low-frequency). For in-domain images, the energy allocation on channels is more compact, with little information of the source images found in low-energy channels. However, for out-of-domain images, more high-frequency components are presented in the spectrum (Fig. 3 (b)), and obvious contours of source images can be observed in reconstructions from low-energy channels, as shown in Fig. 3 (d). This indicates a scattered energy allocation of the pre-trained LIC model on out-of-domain images, which differs from the compact energy allocation on in-domain images.
@@ III-A Motivation
III-B Channel-wise Re-allocation with Adapters
We present the deployment of our method on ELIC [19] in Fig. 4. Two types of adapters, namely Conv-Adapters and LoRA-Adapters, are introduced. To refine the channel-wise decomposition in the transform, we insert Conv-Adapters after non-linear blocks. For entropy estimation, LoRA-Adapters are applied to the entropy parameters network. We highlight the compatibility of the proposed method to mainstream schemes.
Figure 4: Deployment of our method on ELIC [19]. We denote $g_{ep}$ as the entropy parameters network, while the other notations follow the same explanations in [19]. As detailed on the right, Conv-Adapters are inserted serially after non-linear blocks in the transform, while LoRA-Adapters are added to the pre-trained weight matrix $W_{0}$ in $g_{ep}$. $W$ and $b$ are weight and bias of Conv-Adapter, respectively, while $A$ and $B$ are low-rank matrices. Only adapters are trainable.
III-B1 Conv-Adapters
@@
III-B2 LoRA-Adapters
To accommodate the channel-wise re-allocation in the transform, we adopt Low-Rank Adaptation (LoRA) [20; 32] in the entropy model. Mainstream LIC schemes have large variances on the entropy model, resulting in difficulty for a universal adaptation method on entropy estimation. Besides, hidden dimensions in the entropy model are empirically large for better performances, indicating convolution-based adapters will boost the amount of additional parameters. Therefore, we introduce low-rank matrices (LoRA-Adapters) to the common entropy parameters network. Namely, we modify the weight of a pre-trained layer $W_{0}{\in}\mathbb{R}^{c_{out}\times c_{in}\times h\times w}$ by:
@@ III-B2 LoRA-Adapters
III-C Two-Stage Training Strategy
@@
We suggest adapters optimized by a joint objective function suffer from insufficient training under limited samples. Inspired by the two-stage training strategy in [18], we further conduct a second training stage. Concretely, we divide $g_{s}$ into four stacks by upsample operations, as shown in Fig. 5. We freeze all the adapters except the ones in Stack 4, and finetune this small set (denoted as $\Delta\theta$) with hard quantization and distortion only:
@@
IV Experiments
IV-A Settings
IV-A1 Domains and Datasets
We follow Lv et al. [32]; Shen et al. [38] and sort five domains with 150 images each, including (i) pixel-style images [32]; (ii) screen content images from SCID [34], SCI1K [48], SIQAD [47]; (iii) craters images cropped from Lunar Reconnaissance Orbiter Camera (LROC) [36]; (iv) game images cropped from GamingVideoSET [6]; (v) pathology images cropped from BRACS [8]. All crops have the same resolution of $600\times 800$. We select a fixed test set with 100 images for each domain. Besides, we use Kodak to represent the natural image domain.
IV-A2 Models
We test our method on advanced LIC models, namely Cheng2020 [11], ELIC [19], MLIC++ [21] and TCM [30]. We use the official pre-trained weights except ELIC, for which we train ourselves for a lack of official weights.
IV-A3 Implementation
We train adapters with $N=\left\{5,10,25\right\}$ samples, respectively. We split the samples into training and validation set in a proportion of 4:1. We use a patch size of 256 and a batch size of 4. In the first stage, we set learning rate stages $\left\{50,10,7.5,5,2.5,1\right\}\times 10^{-5}$. For different $N$, we further set max epoch =$=\left\{250,500,750\right\}$. In the second stage, we train selected adapters for another max epoch with fixed learning rate $5\times 10^{-4}$.
TABLE I: The number of additional parameters to transmit compared with full-model finetune and DANICE [24], Note that DANICE can not be applied to ELIC.
@@ IV-A3 Implementation
IV-B Adaptation Performance
TABLE II: Computational complexity comparisons of the proposed method and different source models. The results are averaged on Kodak with a single Tesla V100 GPU.

### v2

@@ Abstract
Introduction
With the rapid growth of images in multimedia, image compression has become essential for efficient storage and transmission. Over the past decade, conventional methods including JPEG (Wallace 1991), H.265/HEVC (Sullivan et al. 2012) and H.266/VVC (Bross et al. 2021), have significantly contributed to the real-world applications of image compression. Recently, learned image compression (LIC) has achieved impressive progress in rate-distortion (RD) performance using non-linear transforms and end-to-end optimization. Several studies (He et al. 2022; Jiang and Wang 2023) have even surpassed VTM (the reference software for H.266/VVC) on both PSNR and MS-SSIM, indicating LIC’s potential for future image compression techniques.
@@ Introduction
Despite the encouraging progress, some studies (Shen, Yue, and Yang 2023; Lv et al. 2023) have pointed out the performance degradation of pre-trained LIC models on out-of-domain images (we simplify in/out-of-training-domain as in/out-of-domain for convenience). It is common practice to train and evaluate LIC models on natural image benchmarks (Franzen 1993; Agustsson and Timofte 2017; Asuni and Giachetti 2014; CLIC 2020; Li et al. 2024b). However, practical codecs may well be challenged by images from various domains, many of which have significant domain gaps compared to natural images. For demonstration purposes, we test advanced LIC models including Cheng2020 (Cheng et al. 2020), ELIC (He et al. 2022), TCM (Liu, Sun, and Katto 2023), and MLIC++ (Jiang and Wang 2023) in Fig. 1. While all these models perform comparably or better than the conventional codec VTM on natural images, they exhibit varying degrees of performance drop and generally fall behind VTM on out-of-domain images such as pixel-style art and screen content, highlighting the necessity for efficient adaptation methods for pre-trained LIC models.
@@
In this regard, most existing methods (Shen, Yue, and Yang 2023; Lv et al. 2023; Tsubota, Akutsu, and Aizawa 2023) suggest instance adaptation (IA), which aims to adapt LIC models to a single image at inference. Despite its effectiveness, IA does not perform actual domain adaptation (DA) for pre-trained models, and is time-consuming and computationally expensive due to its per-image online training. On the other hand, Katakol et al. (2021) proposed DA for LIC using limited target samples. They suggested selective finetune, and adapted pre-trained models to a target domain with one-time training and transmission, demonstrating great adaptation efficiency for deployed models. Nevertheless, the proposed method is restricted to GDN structures (Ballé, Laparra, and Simoncelli 2015), and suffers from large additional parameters on recent schemes (He et al. 2022; Liu, Sun, and Katto 2023; Jiang et al. 2023).
@@
Despite the promising applications like codec calibration for batch transmission, DA for LIC remains largely unexplored. Recently, Presta et al. (2024) pre-trained a gate network on multiple domains, enhancing the generalization capacity of LIC models. However, they did not address the problem of adapting LIC models to a specific domain. Therefore, in this paper, we focus on adapting pre-trained LIC models to a target domain. For practicality, we follow Katakol et al. (2021) and study DA for LIC with limited samples. We also highlight the versatility of the DA method to different domains and mainstream LIC schemes.
@@
Building on these insights, we propose a universal few-shot domain adaptation method for LIC using compact adapters, achieving superior performance across various domains and mainstream LIC schemes. Inspired by (Li et al. 2024a), we explore domain gaps in LIC through disturbed channel-wise decomposition. We demonstrate that pre-trained LIC models suffer from scattered channel-wise energy allocation on out-of-domain images. To address this problem, we introduce convolution-based adapters (Conv-Adapters) and low-rank adapters (LoRA-Adapters) to the pre-trained models for channel-wise re-allocation. These adapters are trained with a few target samples, and are plug-and-play for processing specific domains.
@@
Related Work
Learned Image Compression
Learned image compression has demonstrated competitive potential against traditional codecs. Ballé, Laparra, and Simoncelli (2016) proposed the first end-to-end LIC model with non-linear transforms and uniform quantizer, which was further strengthened by hyperprior (Ballé et al. 2018) and context model (Minnen, Ballé, and Toderici 2018). Cheng et al. (2020) first achieved comparable performance with VTM using Gaussian mixture likelihoods. He et al. (2022) proposed uneven channel autoregressive model, demonstrating pleasant performance and complexity. Recently, Liu, Sun, and Katto (2023) designed Transformer-CNN mixture block to incorporate the ability of both structures for LIC. Jiang et al. (2023) introduced multi-reference context to capture different correlations, achieving the state-of-the-art performance (Jiang and Wang 2023).
@@
In the regard of interpretability, early work (Ballé et al. 2018; Cheng et al. 2020) visualized the latent in LIC to evaluate the preciseness of the entropy model. He et al. (2022) sorted latent channels by energy, suggesting low-frequency components concentrated on a few channels while the rest was extremely sparse. Duan et al. (2022) interpreted latent channels as orthogonal transform coefficients, while Li et al. (2024a) suggested frequency decomposition is an intrinsic property for LIC. Nevertheless, these results on latent representation did not fully consider out-of-domain images.
@@
Adaptation for Learned Image Compression
Two kinds of adaptation namely instance adaptation (IA) and domain adaptation (DA) have been studied for LIC. IA (Campos et al. 2019; Van Rozendaal et al. 2021; Tsubota, Akutsu, and Aizawa 2023; Shen, Yue, and Yang 2023; Lv et al. 2023) adapts pre-trained models to a single image rather than a domain using per-image online training. Despite its effectiveness, IA is computationally expensive and time-consuming as shown in Appendix A. Contrarily, Katakol et al. (2021) introduced the problem of DA for LIC, adapting pre-trained LIC models to a target domain with one-time training. They suggested finetuning GDN layers with channel-wise parameters and the entire entropy model, which was restricted by specific structures, and additional parameters due to the complex entropy models. We alleviate this issue by designing compact and universal adapters.
Method
Motivation
Recent studies (Duan et al. 2022; Li et al. 2024a) suggest that the transform in LIC can be interpreted as channel-wise frequency decomposition in a non-linear fashion. We compare latent channels to frequency components of a source image using linear transforms (e.g. FFT), as reconstructions from partial channels or partial frequency components demonstrate obvious similarities in Fig. 2. Motivated by this, we further study the channel-wise decomposition on different domains by reconstructions from specific channels in Fig. 3 (c) and (d). Consistent with previous findings (He et al. 2022; Duan et al. 2022; Li et al. 2024a), decomposed information concentrates on high-energy channels (respectively low-frequency). For in-domain images, the energy allocation on channels is more compact, with little information of the source images found in low-energy channels. However, for out-of-domain images, more high-frequency components are presented in the spectrum (Fig. 3 (b)), and obvious contours of source images can be observed in reconstructions from low-energy channels, as shown in Fig. 3 (d). This indicates a scattered energy allocation of the pre-trained LIC model on out-of-domain images, which differs from the compact energy allocation on in-domain images.
@@
Channel-wise Re-allocation with Adapters
We present the deployment of our method on ELIC (He et al. 2022) in Fig. 4. Two types of adapters, namely Conv-Adapters and LoRA-Adapters, are introduced. To refine the channel-wise decomposition in the transform, we insert Conv-Adapters after non-linear blocks. For entropy estimation, LoRA-Adapters are applied to the entropy parameters network. We highlight the compatibility of the proposed method to mainstream LIC schemes.
Figure 4: Deployment of our method on ELIC (He et al. 2022). We denote $g_{ep}$ as the entropy parameters network, while the other notations follow the same explanations in (He et al. 2022). As detailed on the right, Conv-Adapters are inserted serially after non-linear blocks in the transform, while LoRA-Adapters are added to the pre-trained weight matrix $W_{0}$ in $g_{ep}$. $W$ and $b$ are weight and bias of Conv-Adapter, respectively, while $A$ and $B$ are low-rank matrices. Only adapters are trainable.
Conv-Adapters
@@
LoRA-Adapters
To accommodate the channel-wise re-allocation in the transform, we adopt Low-Rank Adaptation (LoRA) (Hu et al. 2021; Lv et al. 2023) in the entropy model. Mainstream LIC schemes have large variances on the entropy model, resulting in difficulty for a universal adaptation method on entropy estimation. Besides, hidden dimensions in the entropy model are empirically large for better performances, indicating convolution-based adapters will boost the amount of additional parameters. Therefore, we introduce low-rank matrices (LoRA-Adapters) to the common entropy parameters network. Namely, we modify the weight of a pre-trained layer $W_{0}{\in}\mathbb{R}^{c_{out}\times c_{in}\times h\times w}$ by:
@@
Two-Stage Training Strategy
@@
We suggest adapters optimized by a joint objective function suffer from insufficient training under limited samples. Inspired by the two-stage training strategy in (Guo et al. 2021), we further conduct a second training stage. Concretely, we divide $g_{s}$ into four stacks by upsample operations, as shown in Fig. 5. We freeze all the adapters except the ones in Stack 4, and finetune this small set (denoted as $\Delta\theta$) with hard quantization and distortion only:
@@
Experiments
Settings
Domains and Datasets. We follow Lv et al. (2023); Shen, Yue, and Yang (2023) and sort five domains with 150 images each, including (i) pixel-style images (Lv et al. 2023); (ii) screen content images from SCID (Ni et al. 2017), SCI1K (Yang et al. 2021), SIQAD (Yang, Fang, and Lin 2015); (iii) craters images cropped from Lunar Reconnaissance Orbiter Camera (LROC) (Robinson et al. 2010); (iv) game images cropped from GamingVideoSET (Barman et al. 2018); (v) pathology images cropped from BRACS (Brancati et al. 2022). All crops have the same resolution of $600\times 800$. We select a fixed test set with 100 images for each domain. We use Kodak (Franzen 1993) for the natural image domain.
Models. We test our method on advanced LIC models, namely Cheng2020 (Cheng et al. 2020), ELIC (He et al. 2022), MLIC++ (Jiang and Wang 2023) and TCM (Liu, Sun, and Katto 2023). We use the official pre-trained weights except ELIC, for which we train ourselves for a lack of official weights. More details can be found in Appendix F.
Implementation. We train adapters with $N=\left\{5,10,25\right\}$ samples, respectively. We split the samples into training and validation set in a proportion of 4:1. We use a patch size of 256 and a batch size of 4. In the first stage, we set learning rate stages $\left\{50,10,7.5,5,2.5,1\right\}\times 10^{-5}$. For different $N$, we further set max epoch =$=\left\{250,500,750\right\}$. In the second stage, we train selected adapters for another max epoch with fixed learning rate $5\times 10^{-4}$.
@@
Table 1: The number of additional parameters to transmit compared with full-model finetune and DANICE (Katakol et al. 2021), Note that DANICE can not be applied to ELIC.
Adaptation Performance
Comparisons with Traditional Codec
We compare the proposed method with the conventional standard VVC by VTM-22.0 intra coding. We plot the rate-distortion (RD) curves on Kodak (in-domain) and the other five domains in Fig. 6. All pre-trained models perform comparably or better than VTM on Kodak, but most of them fall behind VTM on the other five domains. Our method enhances these models to the level of VTM except on the Pixel domain, where the domain gap is too wide to fully compensate (Lv et al. 2023).
$N$ Baseline Method Pixel Screen-content Craters Game Pathology AVG
  5 Cheng2020 Finetune -23.09% -25.51% 4.35% -10.69% -16.49% -14.29%
DANICE -4.92% -4.48% -7.31% -2.65% -5.62% -5.00%
Ours -21.95% -23.99% -8.55% -8.93% -13.58% -15.40%
ELIC Finetune -15.82% -15.82% -3.26% -2.73% -9.57% -9.44%
DANICE / / / / / /
Ours -17.36% -16.27% -6.07% -4.78% -10.70% -11.04%
DANICE -8.19% -7.78% -10.88% -4.50% -7.14% -7.70%
Ours -27.44% -28.00% -15.25% -11.34% -17.78% -19.96%
ELIC Finetune -20.84% -21.92% -13.06% -9.25% -10.41% -15.10%
DANICE / / / / / /
Ours -21.38% -22.35% -10.64% -6.32% -11.87% -14.51%
DANICE -11.30% -11.59% -13.58% -6.88% -9.83% -10.64%
Ours -27.85% -31.10% -19.12% -14.20% -18.52% -22.16%
ELIC Finetune -24.63% -25.62% -17.15% -15.95% -10.76% -18.82%
DANICE / / / / / /
Ours -23.35% -24.95% -13.77% -13.83% -12.29% -17.64%
TCM Finetune -28.76% -15.53% -13.16% -12.03% -10.21% -15.94%
DANICE -4.12% -4.81% -8.52% 2.30% -4.34% -3.90%
Ours -27.82% -13.81% -8.96% -14.86% -10.71% -15.24%
MLIC++ Finetune -12.97% -9.78% -11.82% -10.61% -9.35% -10.91%
DANICE -0.61% -0.79% -7.15% -0.01% -4.90% -2.69%
Ours -14.63% -12.09% -8.50% -9.70% -10.27% -11.04%
Table 2: Detailed BD-rate ($\downarrow$) using different numbers of target samples ($N$) for domain adaptation. We compare our method with full-model finetune and DANICE. The anchors are corresponding pre-trained models. Note that DANICE can not be applied to ELIC as GDN is not used. Please refer to the Appendix B for more results when $N=\left\{5,10\right\}$.
Stage 1 Stage 2 Pixel Screen-content Craters Game Pathology AVG
Conv LoRA Stack 1234 Stack 4
-19.38% -23.73% -10.84% -12.73% -10.30% -15.40%
-20.40% -24.28% -12.57% -13.40% -11.50% -16.43%
-19.92% -22.46% -11.95% -10.68% -10.95% -15.19%
-23.35% -24.95% -13.77% -13.83% -12.29% -17.64%
Table 3: Ablation studies conducted on ELIC adapted with 25 samples. The anchor of BD-rate ($\downarrow$) is pre-trained ELIC models.
Comparisons with Domain Adaptation Method
To evaluate the efficiency, we also compare with existing DA method DANICE (Katakol et al. 2021) and full-model finetune by RD performance and complexity. We demonstrate detailed adaptation results under different $N$ by computing BD-rate (Bjontegaard 2001) in Table 2. It is noted that DANICE is restricted to GDN structures (Ballé, Laparra, and Simoncelli 2015), making it inapplicable to baselines like ELIC. Additionally, DANICE fails to achieve domain adaptation in some conditions, such as TCM and MLIC++ on the Game domain, indicating a limited versatility. Our method, however, achieves excellent adaptation performance in all conditions, performing comparably with full-model finetune and significantly outperforming DANICE.
In the regard of complexity, we highlight the compactness of adapters in Table 1. The proposed method only needs to transmit fewer than $2\%$ of the total parameters, while much more parameters need considering in DANICE and finetune due to the complex entropy model. Besides, the proposed method costs only a tiny increase on the FLOPs and the decoding time, while enjoys a faster training speed than finetune as suggested in Appendix C.
Ablation Studies
We conduct ablation studies to evaluate the contributions of Conv-Adapters, LoRA-Adapters and the two-stage training strategy. Using pre-trained ELIC as the baseline, we perform the proposed adaptation with 25 samples on the five domains. The results are shown in Table 3. Conv-Adapters achieve an average reduction of 15.40% in BD-rate compared to pre-trained models, while LoRA-Adapters and the two-stage training strategy each contributes an additional 1% BD-rate reduction. We also compare our second stage strategy with finetuning all Conv-Adapters in the decoder and display the superiority of our method.
Given that Conv-Adapter, which involves linear re-allocation on channels, has the most significant impact on adaptation performance, we carry out further explorations on its structure and deployment in Appendix D, and demonstrate the efficiency of the proposed method on both adaptation performance and additional parameters.
Analyses on Channel-wise Re-allocation
Variation on Channels
We use the output latent of the analysis transform, and study the distribution of logarithmic energy and entropy on channels following (He et al. 2022) in Fig. 8. The results are averaged on the Pixel domain. We find that both energy and entropy on specific channels drop after adapting with adapters. Especially, the variation is more apparent on low-energy channels, while high-energy channels remain almost unchanged. This suggests that the energy becomes more concentrated to high-energy channels on out-of-domain images.
Reconstructions from Low-Energy Channels
To further investigate the variation on channel-wise decomposition, we compare the reconstructions from low-energy channels before and after adaptation. Concretely, we keep specific channels unchanged and mask the rest into zeros, reconstruct the modified latent, then subtract the reconstruction of all-zero one. We sort the channels by energy, and display the reconstructions from specific ranges of channels after $200th$ in Fig. 7, which is similar to Fig. 3. We find weaker details of inputs can be observed in reconstructions from the model adapted with our method, suggesting high-frequency components have been better moved out of the low-energy channels, which is more in line with the compact energy allocation in processing in-domain images. Thus, we suggest the pre-trained frequency decomposition on high-frequency components is strengthened by our method. More reconstructions of images from different domains are shown in Appendix E.
Conclusion
Acknowledgments
This work was supported by the Natural Science Foundation of China under Grant 62036005. We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC.
Appendix
Appendix A Comparisons with Instance Adaptation
In this section, we present our proposed domain adaptation (DA) method for learned image compression (LIC), comparing it to existing instance adaptation (IA) methods (Campos et al. 2019; Van Rozendaal et al. 2021; Tsubota, Akutsu, and Aizawa 2023; Shen, Yue, and Yang 2023; Lv et al. 2023). IA adapts a pre-trained LIC model to individual images through per-image online training during inference. In contrast, DA involves an one-time training process on multiple images, adapting the model to a target domain before inference. Consequently, we suggest that our proposed DA method for LIC complements existing IA methods. The compatibility between these methods is illustrated in Fig. 9.
Specifically, we explore the combination of our proposed method with one of the latest IA methods (Lv et al. 2023). We apply both techniques to the Pixel and Screen-content domain. For alignment, we use the pre-trained Cheng2020 model (Cheng et al. 2020) as the baseline. Our DA method is conducted with 25 target samples, followed by implementing IA for each image on the adapted model. The rate-distortion (RD) curves, compared with VTM-22.0 intra coding, the pre-trained Cheng2020, and each technique individually, are shown in Fig. 9. Note that some curves differ slightly due to the different padding method adopted by Lv et al. (2023) in their released code; we respect their implementation in Fig. 9. When combining both techniques, the RD performance is further improved, indicating that our method complements IA by introducing additional domain information from a limited number of samples.
For encoding time, IA relies on expensive per-image online training, whereas DA is more efficient due to its one-time training on multiple samples. We display the encoding time for processing different numbers of test images in Fig. 10. For DA, a fixed one-time training is required to adapt the pre-trained LIC model to the target domain. Adapting the pre-trained Cheng2020 model with 25 samples takes less than half an hour on a single Tesla V100 GPU. Once adapted, the model achieves the same encoding speed as the source model. In contrast, IA requires time-consuming online training for every test image, resulting in a significantly longer encoding time compared to the source model. Therefore, DA is considerably more efficient than IA when encoding a batch of target-domain images.
Figure 9: Compatibility between our method and representative instance adaptation method (Lv et al. 2023) (abbreviated as DA and IA, respectively).
Figure 10: Encoding time of different methods when applied to different numbers of test images from the Pixel domain, using a single Tesla V100 GPU. The proposed few-shot domain adaptation method is conducted on 25 target samples.
Appendix B Detailed Domain Adaptation Performance
We present the detailed domain adaptation performance in Table 6, which is a superset of Table 2. We compare our proposed method with the existing domain adaptation method DANICE (Katakol et al. 2021), and full-model finetune using different numbers of target-domain samples. To evaluate the performance, we compute the BD-rate (Bjontegaard 2001) against the source model. It should be noted that DANICE is restricted to GDN structures (Ballé, Laparra, and Simoncelli 2015), making it inapplicable to baselines like ELIC (He et al. 2022). Our method outperforms DANICE in all conditions and performs comparably with full-model fine-tuning.